### v1

@@ 7 Discussion
This paper presents a systematic study on long-text factuality evaluation by LLMs. Nonetheless, limitations remain in dataset representativeness, multi-agent efficiency, and communication reliability. Future research can explore the following directions:
Expanding and grounding factuality evaluation datasets. While LongHalluQA extends existing short-text datasets to support long-text evaluation, the data are still derived from curated sources rather than real-world user-generated content. Future research should incorporate diverse, high-risk domains such as biomedical, finance, and law, and integrate practical real-world sources to improve coverage, generalization, and applicability of long-text factuality evaluation.
@@
Mitigating communication hallucinations in multi-agent systems. During multi-agent debates, agents may reach misleading consensus due to confidently incorrect responses, causing erroneous conclusions. Yoffe et al. [67] attribute this issue to communication failures among agents, referred to as communication hallucinations. Future research can focus on: (i) confidence-based weighting of agent responses to downweight unreliable viewpoints; (ii) cross-agent consistency verification using logical reasoning to detect contradictions; (iii) adversarial training that injects incorrect viewpoints to test and enhance system robustness. Progress in these areas would strengthen the reliability of multi-agent systems in complex long-text evaluation tasks.

### v2

@@ 7 Discussion
This paper presents a systematic study on long-form factuality evaluation by LLMs. Nonetheless, limitations remain in dataset representativeness, multi-agent efficiency, and communication reliability. Future research can explore the following directions:
Expanding and grounding factuality evaluation datasets. While LongHalluQA extends existing short-text datasets to support long-form evaluation, the data are still derived from curated sources rather than real-world user-generated content. Future research should incorporate diverse, high-risk domains such as biomedical, finance, and law, and integrate practical real-world sources to improve coverage, generalization, and applicability of long-form factuality evaluation.
@@
Mitigating communication hallucinations in multi-agent systems. During multi-agent debates, agents may reach misleading consensus due to confidently incorrect responses, causing erroneous conclusions. Yoffe et al. [67] attribute this issue to communication failures among agents, referred to as communication hallucinations. Future research can focus on: (i) confidence-based weighting of agent responses to downweight unreliable viewpoints; (ii) cross-agent consistency verification using logical reasoning to detect contradictions; (iii) adversarial training that injects incorrect viewpoints to test and enhance system robustness. Progress in these areas would strengthen the reliability of multi-agent systems in complex long-form evaluation tasks.